OpenStudy (anonymous):

v=3k divided by t for t ... would i write v * t = t* 3k divided by t ?

OpenStudy (anonymous):

do you mean the formula is v=3k/t and you have to solve for t?

OpenStudy (anonymous):

yes and yes. i believe i have to get t by itself

OpenStudy (anonymous):

ok well if you want t what should you do with the 3k to get it away from t?

OpenStudy (anonymous):

multiply ?

OpenStudy (anonymous):

multiply it by what?

OpenStudy (anonymous):

multiply 3k by t ?

OpenStudy (anonymous):

no ok think of it as what can you do to 3k to take it away from t and on to the other side?

OpenStudy (anonymous):

subtract t from both sides. so it woould b t-v = 3k - t ?

OpenStudy (anonymous):

ok think of it as... have you been taught how to multiply by a reciprocal?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok so your problem is v=3k/t... you want t by itself. the reciprocal of a number, fraction, etc is the opposite of itself... so like take just 3k and put it over one like this 3k/1 the reciprocal of that is 1/3k.. did you get it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so do you understand how to get 3k away from t now?

OpenStudy (raffle_snaffle):

Rose just making a suggestion. Draw him the procedure on how to eliminate t from one side.

OpenStudy (anonymous):

multiply 3k by t ? :/

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

ok i'll think i'll do that thanks raffle!

OpenStudy (anonymous):

troyanna let me draw it for you ok

OpenStudy (zzr0ck3r):

\[v=\frac{3k}{t}\]multiply both sides by t \[t*v = t*\frac{3k}{t}\\t*v = \cancel{t}*\frac{3k}{\cancel{t}\\tv=3k}\]

OpenStudy (zzr0ck3r):

you with me?

OpenStudy (anonymous):



OpenStudy (anonymous):

yes i had it similar to that in the question

OpenStudy (anonymous):

sorry 2 words got cut out

OpenStudy (zzr0ck3r):

now divide by v to isolate the t \[tv=3k \\\frac{tv}{v}=\frac{3k}{v}\\\frac{t*\cancel{v}}{\cancel{v}}=\frac{3k}{v}\\t=\frac{3k}{v}\]

OpenStudy (zzr0ck3r):

look at it like this \[v =\frac{3k}{t}\\\text{switch the t and the v}\\t=\frac{3k}{v}\]

OpenStudy (zzr0ck3r):

@Rose6 I think you have a mistake

OpenStudy (anonymous):

ok what do you think i made a mistake on? i want to know so next time i can get it right

OpenStudy (anonymous):

wow lol i see what you mean now

OpenStudy (anonymous):

i think we had the same concept i just shortened the way i did it

OpenStudy (anonymous):

did this help you any? Troyanna

OpenStudy (zzr0ck3r):

but you got as the answer \[t=\frac{v}{3k}\]

OpenStudy (zzr0ck3r):

it should be \[t=\frac{3k}{v}\]

OpenStudy (anonymous):

yes. & thank you for the picture

OpenStudy (anonymous):

yes because when you have the problem it was asking to isolate t... well the only way to isolate t is to get rid of 3k by multiplying by its reciprocal on both. then you are done... and no problem Troyanna just glad to help

OpenStudy (anonymous):

*sorry i meant both sides

OpenStudy (zzr0ck3r):

you have a mistake, because the answer is wrong \[v=\frac{3k}{t}\rightarrow t=\frac{3k}{v}\\v=\frac{3k}{t}\cancel{\rightarrow}t=\frac{v}{3k}\]

OpenStudy (anonymous):

i don't get it.... i isolated t by multiplying the reciprocal what is the matter with the end result.... is it because v is at the top and 3k is the denominator?

OpenStudy (anonymous):

the directions were to isolate t from the formula

OpenStudy (zzr0ck3r):

\[\frac{1}{3k}v=\frac{1}{3k}*\frac{3k}{t}=\frac{1}{t}\]

OpenStudy (zzr0ck3r):

understand?

OpenStudy (anonymous):

no i meant to multiply 3k by its reciprocal like this v=3k/t 1/3k*v=3k/t*1/3k 3k cancels on the right side and you are left with v/3k=t

OpenStudy (zzr0ck3r):

lol, I cant do this all day, but its not the right answer.

OpenStudy (anonymous):

v would be at the top because were you to put it over one it would be v/1.... ok thanks for trying to help me though.... i guess i was just taught a bit differently

OpenStudy (zzr0ck3r):

its not about being taught differently, you are not getting the right answer. so if this is the case, you were taught wrong. So you should figure out where the mistake it in your work, because you are getting the wrong answer

OpenStudy (anonymous):

i looked at the website i'm not sure what to tell you i see where you got your answer... ok

OpenStudy (zzr0ck3r):

im not trying to harp on you, i just want you to understand

OpenStudy (anonymous):

no its ok i understand... i would want for some one to tell me if i have the wrong answer however i am the type of person that would always check 500 times on my math so if that is the case then i will look into it Thanks!

OpenStudy (zzr0ck3r):

\[v=\frac{3k}{t}\]multiply both sies by \[\frac{t}{v}\\then\\\frac{t}{v}*v = \frac{t}{v}*\frac{3k}{t}\\\frac{t}{\cancel{v}}\cancel{v}=\frac{\cancel{t}}{v}*\frac{3k}{\cancel{t}}\\t = \frac{3k}{v}\]

OpenStudy (anonymous):

ok thanks!